THE PURE ROTATIONAL SPECTRUM OF ZnCl (X^{2}Σ^{+})

\underline{EMILY D. TENENBAUM}, MICHAEL A. FLORY, ROBIN L. PULLIAM, and LUCY M. ZIURYS, Steward Observatory, University of Arizona, 933 N. Cherry Avenue, Tucson, AZ 85721.

The radical ZnCl has been studied with high-resolution millimeter-wave spectroscopy in its ^{2}Σ^{+} ground state. Pure rotational spectra of ^{64}Zn^{35}Cl, ^{64}Zn^{37}Cl, and ^{66}Zn^{35}Cl were measured in the ground and excited vibrational states (v=1,2), and measurements of ^{67}Zn^{35}Cl, ^{66}Zn^{37}Cl, and ^{68}Zn^{35}Cl were recorded in the ground state. Each rotational transition was found to be split into doublets by spin-rotation interactions, and for ^{67}Zn^{35}Cl, hyperfine splittings due to the nuclear spin of ^{67}Zn (I=5/2) were also observed. Rotational, fine structure, and hyperfine constants have been determined from these data, and equilibrium parameters calculated. The equilibrium bond length of ^{64}Zn^{35}Cl was found to be 2.13003(57)Å, in good agreement with recent theoretical predictions. Interpretation of the hyperfine constants indicates that the ZnCl bond is mostly ionic.